# CarPlanner: Consistent Auto-regressive Trajectory Planning for Large-scale Reinforcement Learning in Autonomous Driving

### v2

@@ Abstract
footnotetext: Project lead. ${}^{\textrm{\Letter}}$ Corresponding author.
@@ 1 Introduction
Most existing researches [3, 34, 13, 19] employ imitation learning (IL) to align planned trajectories with those of human experts. However, this approach suffers from distribution shift [33] and causal confusion [10]. Reinforcement learning (RL) offers a potential solution, addressing these challenges and providing richer supervision through reward functions. Although RL shows effectiveness in domains such as games [39], robotics [22], and language models [28], it still struggles with training inefficiencies and performance issues in the large-scale driving task. To the extent of our knowledge, no RL methods have yet achieved competitive results on large-scale open datasets such as nuPlan [2], which features diverse real-world scenarios.
@@ 2.1 Imitation-based Planning
The use of IL [17, 8] to train planners based on human demonstrations has garnered significant interest recently. This approach leverages the driving expertise of experienced drivers who can safely and comfortably navigate a wide range of real-world scenarios, along with the added advantage of easily collectible driving data at scale [11, 18, 2]. Numerous studies [6, 31, 34, 20] have focused on developing innovative networks to enhance open-loop performance in this domain. However, the ultimate challenge of autonomous driving is achieving closed-loop operation, which is evaluated using driving-oriented metrics such as safety, adherence to traffic rules, comfort, and progress. This reveals a significant gap between the training and testing phases of planners. Moreover, IL is particularly vulnerable to issues such as distribution shift [33] and causal confusion [10]. The first issue results in suboptimal decisions when the system encounters scenarios that are not represented in the training data distribution. The second issue arises when networks inadvertently capture incorrect correlations and develop shortcuts based on input information, primarily due to the reliance on imitation loss from expert demonstrations. Despite efforts in several studies [1, 5, 47, 4] to address these challenges, the gap between training and testing remains substantial.
@@ 2.2 RL in Autonomous Driving
In the field of autonomous driving, RL has demonstrated effectiveness in addressing specific scenarios such as highway driving [25, 44], lane changes [26, 16], and unprotected left turns [25, 27, 45]. Most methods directly learn policies over the control space, which includes throttle, brake, and steering commands. Due to the high frequency of control command execution, the simulation process can be time-consuming, and exploration can be inconsistent [45]. Several works [49, 45] have proposed learning trajectory planners with actions defined as ego-planned trajectories, which temporally extend the exploration space and improve training efficiency. However, a trade-off exists between the trajectory horizon and training performance, as noted in ASAP-RL [45]. Increasing the trajectory horizon results in less reactive behaviors and a reduced amount of data, while a smaller trajectory horizon leads to challenges similar to those encountered in control space. Additionally, these methods typically employ a model-free setting, making them difficult to apply to the complex, diverse real-world scenarios found in large-scale driving datasets. In this paper, we propose adopting a model-based formulation that can facilitate RL training on large-scale datasets. Under this formulation, we aim to overcome the trajectory horizon trade-off by using a world model, which can provide a preview of the world in which our policy can make multi-step decisions during testing.
@@ 3.1 Preliminaries
MDP is used to model sequential decision problems, formulated as a tuple $\mathcal{S},\mathcal{A},P,R,\rho_{0},\gamma,T\right>$. $\mathcal{S}$ is the state space. $\mathcal{A}$ is the action space. $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ $\Delta(\mathcal{X})$ denotes the set of probability distribution over set $\mathcal{X}$. is the state transition probability. $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ denotes the reward function and is bounded. $\rho_{0}\in\Delta(\mathcal{S})$ is the initial state distribution. $T$ is the time horizon and $\gamma$ is the discount factor of future rewards. The state-action sequence is defined as $\tau=(\boldsymbol{s}_{0},a_{0},\boldsymbol{s}_{1},a_{1},\dots,\boldsymbol{s}_{T})$, where $\boldsymbol{s}_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$ are the state and action at time step $t$. The objective of RL is to maximize the expected return:
$\begin{split}\max_{\pi}\mathbb{E}_{\boldsymbol{s}_{t}\sim{P},a_{t}\sim\pi}[\sum_{t=0}^{T}\gamma^{t}R(\boldsymbol{s}_{t},a_{t})].\end{split}$ (1)
@@ 3.2 Problem Formulation
$\begin{split}&P(\boldsymbol{s}_{0},\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{T})=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}P(\boldsymbol{s}_{t+1}|\boldsymbol{s}_{t})\\ &=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}P(s^{0}_{t+1},s^{1:N}_{t+1}|\boldsymbol{s}_{t})\\ &=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}\underbrace{\pi(a_{t}|\boldsymbol{s}_{t})}_{\text{Policy}}\underbrace{P(s^{1:N}_{t+1}|\boldsymbol{s}_{t})}_{\text{Transition Model}}.\\[-4.26773pt] \end{split}$ (3)
@@ 3.3 Planner Architecture
The framework of our proposed CarPlanner is illustrated in Fig. 2, comprising four key components: 1) the non-reactive transition model, 2) the mode selector, 3) the trajectory generator, and 4) the rule-augmented selector. Our planner operates within a generation-selection framework. Given an initial state $s_{0}$ and all possible $N_{\text{mode}}$ modes, the trajectory selector evaluates and assigns scores to each mode. The trajectory generator then produces $N_{\text{mode}}$ trajectories that correspond to their respective modes.
For trajectory generator, the initial state $s_{0}$ is replicated $N_{\text{mode}}$ times, each associated with one of the $N_{\text{mode}}$ modes, effectively creating $N_{\text{mode}}$ parallel worlds. The policy is executed within these simulated environments. During the policy rollout, a trajectory predictor acts as the state transition model, generating future positions of traffic agents across all time horizons. The details are as follows:
@@ 3.3.2 Mode Selector
This module takes $s_{0}$ and longitudinal-lateral decomposed mode information as input and outputs the probability of each mode.
Route-speed decomposed mode. To capture the longitudinal behaviors, we generate $N_{\text{lon}}$ modes that represent the average speed of the trajectory associated with each mode. Each longitudinal mode $c_{\text{lon},j}$ is defined as a scalar value of $\frac{j}{N_{\text{lon}}}$, repeated across a dimension $D$. As a result, the dimensionality of the longitudinal modes is $N_{\text{lon}}\times D$. For lateral behaviors, we identify $N_{\text{lat}}$ possible routes from the map using a graph search algorithm. These routes correspond to the lanes available for the ego vehicle. The dimensionality of these routes is $N_{\text{lat}}\times N_{r}\times D_{m}$. To extract meaningful representations, we employ another PointNet to aggregate the features of the $N_{r}$ points along each route, producing a lateral mode with a dimension of $N_{\text{lat}}\times D$. To create a comprehensive mode representation $\boldsymbol{c}$, we combine the lateral and longitudinal modes, resulting in a combined dimension of $N_{\text{lat}}\times N_{\text{lon}}\times 2D$. To align this mode information with other feature dimensions, we pass it through a linear layer, mapping it back to $N_{\text{lat}}\times N_{\text{lon}}\times D$. $N_{\text{mode}}=N_{\text{lat}}N_{\text{lon}}$.
@@ 3.3.3 Trajectory Generator
This module operates in an auto-regressive manner, recurrently decoding the next pose of the ego vehicle, $a_{t}$, given the current state, $\boldsymbol{s}_{t}$, and consistent mode information, $\boldsymbol{c}$.
Invariant-view module (IVM). Before feeding the mode and state into the network, we preprocess them to eliminate time information. For the map and agent information in state $\boldsymbol{s}_{t}$, we select the $K$ nearest neighbors (KNN) [29] to the ego current pose and only feed these into the policy. $K$ is set to the half of map and agent elements respectively. Regarding the routes that capture lateral behaviors, we filter out the segments where the point closest to the current pose of the ego vehicle is the starting point, retaining $K_{r}$ points. In this case, $K_{r}$ is set to a quarter of $N_{r}$ points in one route. Finally, we transform the routes, agent, and map poses into the coordinate frame of the ego vehicle at the current time step $t$. We subtract the historical time steps $t-H:t$ from the current time step $t$, yielding time steps in range $-H:0$.
Query-based Transformer decoder. We employ the same backbone network architecture as the mode selector, but with different query dimensions. Due to the IVM and the fact that different modes yield distinct states, the map and agent information cannot be shared among modes. As a result, we fuse information for each individual mode. Specifically, the query dimension is $1\times D$, while the dimensions of the keys and values are $(N+N_{m})\times D$. The output feature dimension remains $1\times D$. It is important to highlight that the Transformer decoder can process information from multiple modes in parallel, eliminating the need to handle each mode sequentially using a for loop.
@@ 3.3.4 Rule-augmented Selector
This module first introduces a rule-based selector that takes the initial state $s_{0}$, the multi-modal ego-planned trajectories, and the predicted future trajectories of agents as input. It calculates driving-oriented metrics such as safety, progress, comfort, and others. A comprehensive score is obtained by the weighted sum of the rule-based scores and the mode scores provided by the mode selector. The ego-planned trajectory with the highest score is selected as the output of the trajectory planner.
@@ 3.4 Training
Mode assignment. The positive lateral mode is determined by the endpoint of the ground-truth trajectory. The longitudinal distance from the starting position to this endpoint is divided into $N_{\text{lon}}$ intervals, with the positive longitudinal mode corresponding to the relevant distance interval.
Loss function. For the selector, we use cross-entropy loss that is the negative log-likelihood of the positive mode and a side task that regresses the ego ground-truth trajectory. For the generator, we use PPO [36] loss that consists of three parts: policy improvement, value estimation, and entropy. The full description can be found in the supplementary.
@@
Mode dropout. To prevent over-reliance on mode or route information due to Transformers’ residual connections, we implement a mode dropout module during training that randomly masks the route to mitigate this issue.
@@ 4.2 Comparison with SOTAs
SOTAs. We categorize the methods into Rule, IL, and RL based on the type of trajectory generator. (1) PDM [9] wins the nuPlan challenge 2023, its IL-based and rule-based variants are denoted as PDM-Closed and PDM-Open, respectively. PDM-Closed follows the generation-selection framework where IDM is used to generate multiple candidate trajectories and rule-based selector considering safety, progress, and comfort is used to select the best trajectory. (2) PLUTO [4] also obeys the generation-selection framework and uses contrastive IL to incorporate various data augmentation techniques and trains the generator. (3) Gen-Drive [21] is a concurrent work that follows a pretrain-finetune pipeline where IL is used to pretrain a diffusion-based planner and RL is used to finetune the denoising process based on a reward model trained by AI preference.
@@ 4.5 Qualitative Results
We provide qualitative results as shown in Fig. 3. In this scenario, ego vehicle is required to execute a right turn while navigating around pedestrians. In this case, Our method shows a smooth, efficient performance. From $t_{\text{sim}}=0s$ to $t_{\text{sim}}=9s$, all methods wait for the pedestrians to cross the road. At $t_{\text{sim}}=10s$, an unexpected pedestrian goes back and prepares to re-cross the road. PDM-Closed is unaware of this situation and takes an emergency stop but still intersects with this pedestrian. In contrast, our IL variant displays an awareness of the pedestrian’s movements and consequently conducts a braking maneuver. However, it still remains close to the pedestrian. Our RL method avoids this hazard by starting up early up to $t_{\text{sim}}=9s$ and achieves the highest progress and safety metrics.
@@ 5 Conclusion
Limitations and future work. RL needs delicate design and is prone to input representation. Besides, RL can overfit its training environment and suffer from performance drop in unseen environments [24]. Future work aims to develop robust RL algorithms to tackle these limitations.

### v3

@@ Abstract
11footnotetext: Project lead. ${}^{\textrm{\Letter}}$ Corresponding author.
@@ 1 Introduction
Most existing researches [29, 11, 16] employ imitation learning (IL) to align planned trajectories with those of human experts. However, this approach suffers from distribution shift [28] and causal confusion [8]. Reinforcement learning (RL) offers a potential solution, addressing these challenges and providing richer supervision through reward functions. Although RL shows effectiveness in domains such as games [34], robotics [19], and language models [25], it still struggles with training inefficiencies and performance issues in the large-scale driving task. To the extent of our knowledge, no RL methods have yet achieved competitive results on large-scale open datasets such as nuPlan [2], which features diverse real-world scenarios.
@@ 2.1 Imitation-based Planning
The use of IL to train planners based on human demonstrations has garnered significant interest recently. This approach leverages the driving expertise of experienced drivers who can safely and comfortably navigate a wide range of real-world scenarios, along with the added advantage of easily collectible driving data at scale [9, 15, 2]. Numerous studies [5, 29, 17] have focused on developing innovative networks to enhance open-loop performance in this domain. However, the ultimate challenge of autonomous driving is achieving closed-loop operation, which is evaluated using driving-oriented metrics such as safety, adherence to traffic rules, comfort, and progress. This reveals a significant gap between the training and testing phases of planners. Moreover, IL is particularly vulnerable to issues such as distribution shift [28] and causal confusion [8]. The first issue results in suboptimal decisions when the system encounters scenarios that are not represented in the training data distribution. The second issue arises when networks inadvertently capture incorrect correlations and develop shortcuts based on input information, primarily due to the reliance on imitation loss from expert demonstrations. Despite efforts in several studies [1, 4, 42, 3] to address these challenges, the gap between training and testing remains substantial.
@@ 2.2 RL in Autonomous Driving
In the field of autonomous driving, RL has demonstrated its effectiveness in addressing specific scenarios such as highway driving [22, 39], lane changes [23, 14], and unprotected left turns [22, 40]. Most methods directly learn policies over the control space, which includes throttle, brake, and steering commands. Due to the high frequency of control command execution, the simulation process can be time-consuming, and exploration can be inconsistent [40]. Several works [44, 40] have proposed learning trajectory planners with actions defined as ego-planned trajectories, which temporally extend the exploration space and improve training efficiency. However, a trade-off exists between the trajectory horizon and training performance, as noted in ASAP-RL [40]. Increasing the trajectory horizon results in less reactive behaviors and a reduced amount of data, while a smaller trajectory horizon leads to challenges similar to those encountered in control space. Additionally, these methods typically employ a model-free setting, making them difficult to apply to the complex, diverse real-world scenarios found in large-scale driving datasets. In this paper, we propose adopting a model-based formulation that can facilitate RL training on large-scale datasets. Under this formulation, we aim to overcome the trajectory horizon trade-off by using a transition model, which can provide a preview of the world in which our policy can make multi-step decisions during testing.
@@ 3.1 Preliminaries
MDP is used to model sequential decision problems, formulated as a tuple $\mathcal{S},\mathcal{A},P_{\tau},R,\rho_{0},\gamma,T\right>$. $\mathcal{S}$ is the state space. $\mathcal{A}$ is the action space. $P_{\tau}:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ $\Delta(\mathcal{X})$ denotes the set of probability distribution over set $\mathcal{X}$. is the state transition probability. $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ denotes the reward function and is bounded. $\rho_{0}\in\Delta(\mathcal{S})$ is the initial state distribution. $T$ is the time horizon and $\gamma$ is the discount factor of future rewards. The state-action sequence is defined as $\tau=(\boldsymbol{s}_{0},a_{0},\boldsymbol{s}_{1},a_{1},\dots,\boldsymbol{s}_{T})$, where $\boldsymbol{s}_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$ are the state and action at time step $t$. The objective of RL is to maximize the expected return:
$\begin{split}\max_{\pi}\mathbb{E}_{\boldsymbol{s}_{t}\sim{P_{\tau}},a_{t}\sim\pi}[\sum_{t=0}^{T}\gamma^{t}R(\boldsymbol{s}_{t},a_{t})].\end{split}$ (1)
@@ 3.2 Problem Formulation
$\begin{split}&P(\boldsymbol{s}_{0},\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{T})=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}P(\boldsymbol{s}_{t+1}|\boldsymbol{s}_{t})\\ &=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}P(s^{0}_{t+1},s^{1:N}_{t+1}|\boldsymbol{s}_{t})\\ &=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}\underbrace{\pi(a_{t}|\boldsymbol{s}_{t})}_{\text{Policy}}\underbrace{P_{\tau}(s^{1:N}_{t+1}|\boldsymbol{s}_{t})}_{\text{Transition Model}}.\\[-4.26773pt] \end{split}$ (3)
@@ 3.3 Planner Architecture
The framework of our proposed CarPlanner is illustrated in Fig. 2, comprising four key components: 1) the non-reactive transition model, 2) the mode selector, 3) the trajectory generator, and 4) the rule-augmented selector.
Our planner operates within a generation-selection framework. Given an initial state $s_{0}$ and all possible $N_{\text{mode}}$ modes, the trajectory selector evaluates and assigns scores to each mode. The trajectory generator then produces $N_{\text{mode}}$ trajectories that correspond to their respective modes. For trajectory generator, the initial state $s_{0}$ is replicated $N_{\text{mode}}$ times, each associated with one of the $N_{\text{mode}}$ modes, effectively creating $N_{\text{mode}}$ parallel worlds. The policy is executed within these previewed worlds. During the policy rollout, a trajectory predictor acts as the state transition model, generating future poses of traffic agents across all time horizons.
@@ 3.3.2 Mode Selector
This module takes $s_{0}$ and longitudinal-lateral decomposed mode information as input and outputs the probability of each mode. The number of modes $N_{\text{mode}}=N_{\text{lat}}N_{\text{lon}}$.
Route-speed decomposed mode. To capture the longitudinal behaviors, we generate $N_{\text{lon}}$ modes that represent the average speed of the trajectory associated with each mode. Each longitudinal mode $c_{\text{lon},j}$ is defined as a scalar value of $\frac{j}{N_{\text{lon}}}$, repeated across a dimension $D$. As a result, the dimensionality of the longitudinal modes is $N_{\text{lon}}\times D$. For lateral behaviors, we identify $N_{\text{lat}}$ possible routes from the map using a graph search algorithm. These routes correspond to the lanes available for the ego vehicle. The dimensionality of these routes is $N_{\text{lat}}\times N_{r}\times D_{m}$. We employ another PointNet to aggregate the features of the $N_{r}$ points along each route, producing a lateral mode with a dimension of $N_{\text{lat}}\times D$. To create a comprehensive mode representation $\boldsymbol{c}$, we combine the lateral and longitudinal modes, resulting in a combined dimension of $N_{\text{lat}}\times N_{\text{lon}}\times 2D$. To align this mode information with other feature dimensions, we pass it through a linear layer, mapping it back to $N_{\text{lat}}\times N_{\text{lon}}\times D$.
@@ 3.3.3 Trajectory Generator
This module operates in an auto-regressive manner, recurrently decoding the next pose of the ego vehicle $a_{t}$, given the current state $\boldsymbol{s}_{t}$, and consistent mode information $\boldsymbol{c}$.
Invariant-view module (IVM). Before feeding the mode and state into the network, we preprocess them to eliminate time information. For the map and agent information in state $\boldsymbol{s}_{t}$, we select the $K$-nearest neighbors (KNN) to the ego current pose and only feed these into the policy. $K$ is set to the half of map and agent elements respectively. Regarding the routes that capture lateral behaviors, we filter out the segments where the point closest to the current pose of the ego vehicle is the starting point, retaining $K_{r}$ points. In this case, $K_{r}$ is set to a quarter of $N_{r}$ points in one route. Finally, we transform the routes, agent, and map poses into the coordinate frame of the ego vehicle at the current time step $t$. We subtract the historical time steps $t-H:t$ from the current time step $t$, yielding time steps in range $-H:0$.
Query-based Transformer decoder. We employ the same backbone network architecture as the mode selector, but with different query dimensions. Due to the IVM and the fact that different modes yield distinct states, the map and agent information cannot be shared among modes. As a result, we fuse information for each individual mode. Specifically, the query dimension is $1\times D$, while the dimensions of the keys and values are $(N+N_{m})\times D$. The output feature dimension remains $1\times D$. Note that Transformer decoder can process information from multiple modes in parallel, eliminating the need to handle each mode sequentially.
@@ 3.3.4 Rule-augmented Selector
This module is only utilized during inference and takes as input the initial state $s_{0}$, the multi-modal ego-planned trajectories, and the predicted future trajectories of agents. It calculates driving-oriented metrics such as safety, progress, comfort. A comprehensive score is obtained by the weighted sum of rule-based scores and the mode scores provided by the mode selector. The ego-planned trajectory with the highest score is selected as the output of the planner.
@@ 3.4 Training
Mode assignment. For the lateral mode, we assign the route closest to the endpoint of ego ground-truth trajectory as the positive lateral mode. For the longitudinal mode, we partition the longitudinal space into $N_{\text{lon}}$ intervals and assign the interval containing the endpoint of the ground-truth trajectory as the positive longitudinal mode.
@@
Mode dropout. In some cases, there are no available routes for ego to follow. However, since routes serve as queries in Transformer, the absence of a route can lead to unstable or hazardous outputs. To mitigate this issue, we implement a mode dropout module during training that randomly masks routes to prevent over-reliance on this information.
Loss function. For the selector, we use cross-entropy loss that is the negative log-likelihood of the positive mode and a side task that regresses the ego ground-truth trajectory. For the generator, we use PPO [31] loss that consists of three parts: policy improvement, value estimation, and entropy. Full description can be found in supplementary.
@@ 4.2 Comparison with SOTAs
SOTAs. We categorize the methods into Rule, IL, and RL based on the type of trajectory generator. (1) PDM [7] wins the nuPlan challenge 2023, its IL-based and rule-based variants are denoted as PDM-Open and PDM-Closed, respectively. PDM-Closed follows the generation-selection framework where IDM is used to generate multiple candidate trajectories and rule-based selector considering safety, progress, and comfort is used to select the best trajectory. (2) PLUTO [3] also obeys the generation-selection framework and uses contrastive IL to incorporate various data augmentation techniques and trains the generator. (3) Gen-Drive [18] is a concurrent work that follows a pretrain-finetune pipeline where IL is used to pretrain a diffusion-based planner and RL is used to finetune the denoising process based on a reward model trained by AI preference.
@@ 4.5 Qualitative Results
We provide qualitative results as shown in Fig. 3. In this scenario, ego vehicle is required to execute a right turn while navigating around pedestrians. In this case, Our method shows a smooth, efficient performance. From $t_{\text{sim}}=0s$ to $t_{\text{sim}}=9s$, all methods wait for the pedestrians to cross the road. At $t_{\text{sim}}=10s$, an unexpected pedestrian goes back and prepares to re-cross the road. PDM-Closed is unaware of this situation and takes an emergency stop, but it still intersects with this pedestrian. In contrast, our IL variant displays an awareness of the pedestrian’s movements and consequently conducts a braking maneuver. However, it still remains close to the pedestrian. Our RL method avoids this hazard by starting up early up to $t_{\text{sim}}=9s$ and achieves the highest progress and safety metrics.
@@ 5 Conclusion
Limitations and future work. RL needs delicate design and is prone to input representation. RL can overfit its training environment and suffer from performance drop in unseen environments [21]. Our method leverages expert-aided reward design to guide exploration. However, this approach may constrain the full potential of RL, as it inherently relies on expert demonstrations and may hinder the discovery of solutions that surpass human expertise. Future work aims to develop robust RL algorithms capable of overcoming these limitations, enabling autonomous exploration and generalization across diverse environments.